# Regularized least squares learning with heavy-tailed noise is minimax optimal

### v2

@@ C.2 Bennett inequality in Hilbert spaces
C.3 Fuk–Nagaev inequality in Hilbert spaces
The bound in Proposition 5.1 given in the main text is a sharper version of a more general bound given by [27] for general normed spaces, which we state here for completeness.
Proposition C.3 (Fuk–Nagaev inequality, [27], Theorem 3.5.1).
@@ C.3 Fuk–Nagaev inequality in Hilbert spaces
$\mathbb{E}[\xi_{i}]=0,\quad\sum_{i=1}^{n}\mathbb{E}\mathopen{}\mathclose{{}\left[\lVert\xi_{i}\rVert_{\mathcal{X}}^{2}}\right]<B^{2}\quad\text{and}\quad\sum_{i=1}^{n}\mathbb{E}\mathopen{}\mathclose{{}\left[\lVert\xi_{i}\rVert_{\mathcal{X}}^{q}}\right]<A$ (C.1)
for some constants $B^{2}>0$, $A>0$ and $q\geq 3$. Then there exist two universal constants $c_{1}>0$ and $c_{2}>0$ depending only on $q$, such that with $S_{n}:=\sum_{i=1}^{n}\xi_{i}$ we have for every $t>0$ that
@@
Compared to the result above, Proposition 5.1 removes the excess term $\mathbb{E}[\lVert S_{n}\rVert_{\mathcal{X}}]$ by making use of the geometry of the Hilbert space. We now prove Proposition 5.1 based on the proof by [27] and emphasize that this approach can be directly extended to a $D$-smooth Banach space by incorporating arguments by [63] with adjusted constants.
The following statement sharpens [27, Lemma 3.5.1] in the Hilbert space setting.
Lemma C.4.
Let $\xi_{1},\dots\xi_{n}$ be independent random variables taking values in a separable Hilbert space $\mathcal{X}$ such that the conditions (C.1) are satisfied and additionally $\lVert\xi_{i}\rVert_{\mathcal{X}}\leq L$ holds almost surely for some $L>0$ and all $1\leq i\leq n$. Then for all $h>0$, we have
$\mathbb{E}[\cosh(h\lVert S_{n}\rVert_{\mathcal{X}})]\leq\exp\mathopen{}\mathclose{{}\left(\tilde{c}_{0}+\frac{1}{2}h^{2}B^{2}+\tilde{c}_{1}Ah^{q}e^{hL}}\right),$
where $\tilde{c}_{0},\tilde{c}_{1}>0$ are constants depending exclusively on $q$.
@@
$\leq\exp\mathopen{}\mathclose{{}\left(\tilde{c}_{0}+\frac{1}{2}h^{2}B^{2}+\tilde{c}_{1}Ah^{q}e^{hL}}\right),$
@@
where the first inequality is due to [27, Theorem 3.3.4(a)], the second follows from the fact that we have $x\leq\exp(x-1)$ for all $x\in\mathbb{R}$ and the third inequality is shown in the proof of [27, Lemma 3.5.1]. ∎
Proof of Proposition 5.1..
@@
We can now prove Proposition 5.1 by modifying the truncation argument and Chernoff bound given by [27, Section 3.5.2] in combination with Lemma C.4, which sharpens a more general bound that is used in the original proof by [27]. Let us assume that (C.1) holds. For $L>0$, we introduce the truncated random variables
[…]
@@
In contrast to $S_{n}$, the truncated sum $\tilde{S}_{n}$ is not necessarily centered. Here, our version of the proof differs from the proof of [27, Theorem 3.5.1]. The original proof approximates the centered norm $\lVert S_{n}\rVert_{\mathcal{X}}-\mathbb{E}[\lVert S_{n}\rVert_{\mathcal{X}}]$ with the truncated centered norm $\lVert\tilde{S}_{n}\rVert_{\mathcal{X}}-\mathbb{E}[\lVert\tilde{S}_{n}\rVert_{\mathcal{X}}]$, leading to the excess term $\mathbb{E}[\lVert S_{n}\rVert_{\mathcal{X}}]$ in the event for which the tail is bounded. In contrast, we perform the centering directly in the norm: we approximate $\lVert S_{n}\rVert_{\mathcal{X}}$ with $\lVert\tilde{S}_{n}-\mathbb{E}[\tilde{S}_{n}]\rVert_{\mathcal{X}}$.
@@
$\leq\frac{A}{L^{q-1}}.$ (C.2)
We now verify the conditions (C.1) for the centered sum $\tilde{S}_{n}-\mathbb{E}[\tilde{S}_{n}]$. We obviously have
@@
$\leq\frac{A}{L^{q}}+\mathbb{P}[\lVert\tilde{S}_{n}-\mathbb{E}[\tilde{S}_{n}]]\rVert_{\mathcal{X}}\geq tB-A/L^{q-1}]$ (by (C.3))
@@
[…]
$\leq 2\mathopen{}\mathclose{{}\left(\frac{\tilde{A}}{L^{q}}+\exp\mathopen{}\mathclose{{}\left(-htB+\frac{h\tilde{A}}{L^{q-1}}+\tilde{c}_{0}+\frac{1}{2}h^{2}B^{2}+\tilde{c}_{1}\tilde{A}h^{q}e^{hL}}\right)}\right).$
The last expression in the parentheses is minimized over the choices of $L$ and $h$ as shown in detail by [27, pp. 105], giving
$\mathbb{P}\big{[}\,\lVert S_{n}\rVert_{\mathcal{X}}\geq tB\big{]}\leq c_{1}\mathopen{}\mathclose{{}\left(\frac{A}{B^{q}t^{q}}{}+\exp(-c_{2}t^{2})}\right)$ (C.3)
@@
Whenever the $\xi_{i}$ satisfy $\mathbb{E}[\lVert\xi_{i}\rVert_{\mathcal{X}}^{2}]=\sigma^{2}$ and $\mathbb{E}[\lVert\xi_{i}\rVert_{\mathcal{X}}^{q}]=Q$ for some constants $\sigma^{2},Q>0$, we may substitute $B=n\sigma^{2}$ and $A=nQ$ in (C.3). Rearranging proves the bound given in Proposition 5.1. ∎
C.4 Concentration of empirical covariance operators

### v3

@@ D.2 Bennett inequality in Hilbert spaces
D.3 Fuk–Nagaev inequality in Hilbert spaces
The bound in Proposition 5.1 given in the main text is a sharper version of a more general bound given by [Yurinsky1995] for general normed spaces, which we state here for completeness.
The original proof of the result below relies on an inconsistent exponential moment bound, which we address by providing an alternative bound alongside a detailed proof. We were made aware of this fact by Christian Fiedler [christian], who contributed to the alternative proof presented here.
Proposition D.3 (Fuk–Nagaev inequality, [Yurinsky1995], Theorem 3.5.1).
@@ D.3 Fuk–Nagaev inequality in Hilbert spaces
$\mathbb{E}[\xi_{i}]=0,\quad\sum_{i=1}^{n}\mathbb{E}\mathopen{}\mathclose{{\left[\lVert\xi_{i}\rVert_{\mathcal{X}}^{2}}}\right]<B^{2}\quad\text{and}\quad\sum_{i=1}^{n}\mathbb{E}\mathopen{}\mathclose{{\left[\lVert\xi_{i}\rVert_{\mathcal{X}}^{q}}}\right]<A$ (D.1)
for some constants $B^{2}>0$, $A>0$ and $q\in\mathbb{N}$, $q\geq 3$. Then there exist two universal constants $c_{1}>0$ and $c_{2}>0$ depending only on $q$, such that with $S_{n}:=\sum_{i=1}^{n}\xi_{i}$ we have for every $t>0$ that
@@
Compared to the result above, Proposition 5.1 removes the excess term $\mathbb{E}[\lVert S_{n}\rVert_{\mathcal{X}}]$ on the left hand side by making use of the geometry of the Hilbert space norm (see [Pinelis1986Remarks, Pinelis94]).
We now prove Proposition 5.1 and emphasize that it can be directly extended to random variables in $(2,D)$-smooth Banach spaces by incorporating arguments from [Pinelis94] with adjusted constants. We provide the proof in three parts: In Section D.3.1, we give an exponential moment bound. Section D.3.2 contains the core proof of Proposition 5.1 and incorporates the exponential moment bound into a Chernoff bound that is obtained via a truncation argument. Section D.3.3 provides the final optimization of the Chernoff bound.
D.3.1 Exponential moment bound
We now provide a sharpened version of [Yurinsky1995, Lemma 3.5.1] that allows to bound the exponential moments of sums of random variables in Hilbert spaces. The proof of [Yurinsky1995, Lemma 3.5.1] is lacking details and is inconsistent, as it generally seems to require bounded scaling factors $h>0$ in the exponent, but the result is stated for all $h>0$. Our proof holds for all $h>0$ and gives a slightly adjusted bound.
Lemma D.4.
Let $\xi_{1},\dots\xi_{n}$ be independent random variables taking values in a separable Hilbert space $\mathcal{X}$ such that the conditions (D.1) are satisfied and additionally $\lVert\xi_{i}\rVert_{\mathcal{X}}\leq L$ holds almost surely for some $L>0$ and all $1\leq i\leq n$. Then for all $h>0$, we have
$\mathbb{E}[\cosh(h\lVert S_{n}\rVert_{\mathcal{X}})]\leq\exp\mathopen{}\mathclose{{\left(\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}Ah^{q}e^{hL}}}\right),$
where the positive multiplicative constants $\tilde{c}_{2}$ and $\tilde{c}_{q}$ only depend on $q$.
The upper bound of Lemma D.4 differs from [Yurinsky1995, Lemma 3.5.1] in the sense that it removes an additive constant in the exponent. Furthermore, our constant $\tilde{c}_{2}$ of the quadratic term depends on $q$, while $\tilde{c}_{2}=1/2$ for all $q$ in [Yurinsky1995, Lemma 3.5.1], which seems to conflict with unbounded $h>0$.
@@ D.3.1 Exponential moment bound
where the first inequality is due to [Pinelis1986Remarks, Theorem 3], the second inequality follows from the fact that we have $x\leq\exp(x-1)$ for all $x\in\mathbb{R}$. We hence need to show
$\sum_{i=1}^{n}\mathbb{E}[e^{h\lVert\xi_{i}\rVert_{\mathcal{X}}}-1-h\lVert\xi_{i}\rVert_{\mathcal{X}}]\leq\tilde{c}_{1}h^{2}B^{2}+\tilde{c}_{q}Ah^{q}e^{hL}.$
We start by expanding the term on the left hand side as
$\sum_{i=1}^{n}e^{h\lVert\xi_{i}\rVert_{\mathcal{X}}}-1-h\lVert\xi_{i}\rVert_{\mathcal{X}}=\sum_{i=1}^{n}\sum_{j=2}^{\infty}\frac{h^{j}\lVert\xi_{i}\rVert^{j}_{\mathcal{X}}}{j!}.$
The term for the index $j=2$ is immediately bounded as
$\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\frac{h^{2}\lVert\xi_{i}\rVert^{2}_{\mathcal{X}}}{2!}}}\right]\leq\frac{1}{2}h^{2}B^{2}.$
Hence, the claim now reduces to showing
$\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\sum_{j=3}^{\infty}\frac{h^{j}\lVert\xi_{i}\rVert^{j}_{\mathcal{X}}}{j!}}}\right]\lesssim h^{2}B^{2}+Ah^{q}e^{hL}$
with constants only depending on $q$. We split the inner sum over $j$ at the index $q$ and obtain
$\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\sum_{j=3}^{\infty}\frac{h^{j}\lVert\xi_{i}\rVert^{j}_{\mathcal{X}}}{j!}}}\right]=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\sum_{j=3}^{q-1}\frac{h^{j}\lVert\xi_{i}\rVert^{j}_{\mathcal{X}}}{j!}}}\right]+\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\sum_{j=q}^{\infty}\frac{h^{j}\lVert\xi_{i}\rVert^{j}_{\mathcal{X}}}{j!}}}\right].$
We now proceed to bound both terms on the right hand side individually.
Bounding the first summand.
We address the first term as
$\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\sum_{j=3}^{q-1}\frac{h^{j}\lVert\xi_{i}\rVert^{j}_{\mathcal{X}}}{j!}}}\right]$ $\leq\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\sum_{j=3}^{q-1}\frac{\frac{j-2}{q-2}\cdot h^{q}\lVert\xi_{i}\rVert^{q}_{\mathcal{X}}+\frac{q-j}{q-2}\cdot h^{2}\lVert\xi_{i}\rVert^{2}_{\mathcal{X}}}{j!}}}\right]$
$=\sum_{j=3}^{q-1}\frac{1}{j!}\sum_{i=1}^{n}\frac{j-2}{q-2}\cdot h^{q}\mathbb{E}[\lVert\xi_{i}\rVert_{\mathcal{X}}^{q}]+\frac{q-j}{q-2}\cdot h^{2}\mathbb{E}[\lVert\xi_{i}\rVert_{\mathcal{X}}^{2}]$
$=\sum_{j=3}^{q-1}\frac{1}{j!}\mathopen{}\mathclose{{\left(\frac{j-2}{q-2}\cdot h^{q}A+\frac{q-j}{q-2}\cdot h^{2}B^{2}}}\right)$
$\lesssim h^{q}A+h^{2}B^{2}\leq h^{q}Ae^{hL}+h^{2}B^{2},$
where we apply Lemma E.3 with $x=h\lVert\xi_{j}\rVert_{\mathcal{X}}$, $a=1$ and $\rho=j$ to each individual term in the sum and use that $h>0$ and $L>0$ ensure $e^{hL}\geq 1$. Note that the unspecified constants only depend on $q$.
Bounding the second summand.
It remains to bound the second sum. Almost surely, we have
$\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\sum_{j=q}^{\infty}\frac{h^{j}\lVert\xi_{i}\rVert^{j}_{\mathcal{X}}}{j!}}}\right]$ $=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\frac{h^{q}\lVert\xi_{i}\rVert^{q}_{\mathcal{X}}}{q!}+\frac{h^{q+1}\lVert\xi_{i}\rVert^{q+1}_{\mathcal{X}}}{(q+1)!}\cdots}}\right]$
$=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\frac{h^{q}\lVert\xi_{i}\rVert^{q}_{\mathcal{X}}}{q!}\mathopen{}\mathclose{{\left(1+\frac{h\lVert\xi_{i}\rVert_{\mathcal{X}}}{(q+1)}+\frac{h^{2}\lVert\xi_{i}\rVert^{2}_{\mathcal{X}}}{(q+1)(q+2)}\cdots}}\right)}}\right]$
$\leq\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{i=1}^{n}\frac{h^{q}\lVert\xi_{i}\rVert^{q}_{\mathcal{X}}}{q!}\mathopen{}\mathclose{{\left(1+\frac{h^{1}L}{(q+1)}+\frac{h^{2}L^{2}}{(q+1)(q+2)}\cdots}}\right)}}\right]$
$\leq\frac{h^{q}}{q!}e^{hL}\sum_{i=1}^{n}\mathbb{E}\mathopen{}\mathclose{{\left[\lVert\xi_{i}\rVert_{\mathcal{X}}^{q}}}\right]\leq\frac{h^{q}}{q!}e^{hL}A,$
proving the claim. ∎
D.3.2 Core proof of Proposition 5.1
@@
We can now prove Proposition 5.1 by modifying the truncation argument and Chernoff bound given by [Yurinsky1995, Section 3.5.2] in combination with Lemma D.4. Let us assume that (D.1) holds. For $L>0$, we introduce the truncated random variables
$\tilde{\xi}_{i}:=\xi_{i}\,\mathds{1}_{\,[\|\xi_{i}\|_{\mathcal{X}}\leq L\,]},\qquad\tilde{S}_{n}:=\sum_{i=1}^{n}\tilde{\xi}_{i}.$
@@
In contrast to $S_{n}$, the truncated sum $\tilde{S}_{n}$ is not necessarily centered. Here, our version of the proof differs from the proof of [Yurinsky1995, Theorem 3.5.1]. The original proof approximates the centered norm $\lVert S_{n}\rVert_{\mathcal{X}}-\mathbb{E}[\lVert S_{n}\rVert_{\mathcal{X}}]$ with the truncated centered norm $\lVert\tilde{S}_{n}\rVert_{\mathcal{X}}-\mathbb{E}[\lVert\tilde{S}_{n}\rVert_{\mathcal{X}}]$, leading to the excess term $\mathbb{E}[\lVert S_{n}\rVert_{\mathcal{X}}]$ in the event for which the tail is bounded. In contrast, we perform the centering directly in the norm: we approximate $\lVert S_{n}\rVert_{\mathcal{X}}$ with $\lVert\tilde{S}_{n}-\mathbb{E}[\tilde{S}_{n}]\rVert_{\mathcal{X}}$.
@@
$\leq\frac{A}{L^{q-1}}.$ (D.2)
We now verify the conditions (D.1) for the centered sum $\tilde{S}_{n}-\mathbb{E}[\tilde{S}_{n}]$. We obviously have
@@
$\leq\frac{A}{L^{q}}+\mathbb{P}[\lVert\tilde{S}_{n}-\mathbb{E}[\tilde{S}_{n}]]\rVert_{\mathcal{X}}\geq tB-A/L^{q-1}]$ (by (D.3.2))
@@
$\leq\frac{A}{L^{q}}+\exp\mathopen{}\mathclose{{\left(-htB+\frac{hA}{L^{q-1}}}}\right)\,2\mathbb{E}[\cosh(h\lVert\tilde{S}_{n}-\mathbb{E}[\tilde{S}_{n}]\rVert_{\mathcal{X}})]$ ($\cosh(x)=(e^{x}+e^{-x})/2$)
@@
$\leq 2\mathopen{}\mathclose{{\left(\frac{\tilde{A}}{L^{q}}+\exp\mathopen{}\mathclose{{\left(-htB+\frac{h\tilde{A}}{L^{q-1}}+\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}\tilde{A}h^{q}e^{hL}}}\right)}}\right).$
We derive an upper bound of the last expression in the parentheses over all choices of $L>0$ and $h>0$ as shown in detail in Section D.3.3, giving
$\mathbb{P}\big[\,\lVert S_{n}\rVert_{\mathcal{X}}\geq tB\big]\leq c_{1}\mathopen{}\mathclose{{\left(\frac{\tilde{A}}{B^{q}t^{q}}+\exp(-c_{2}t^{2})}}\right)$ (D.3)
@@
Whenever the $\xi_{i}$ satisfy $\mathbb{E}[\lVert\xi_{i}\rVert_{\mathcal{X}}^{2}]=\sigma^{2}$ and $\mathbb{E}[\lVert\xi_{i}\rVert_{\mathcal{X}}^{q}]=Q$ for some constants $\sigma^{2},Q>0$, we may substitute $B=n\sigma^{2}$ and $A=nQ$ in (D.3). Rearranging proves the bound given in Proposition 5.1.
D.3.3 Chernoff bound optimization
The final bound we want to obtain is of the form
$\min_{h,L\geq 0}\frac{\tilde{A}}{L^{q}}+\exp\mathopen{}\mathclose{{\left(-htB+\frac{h\tilde{A}}{L^{q-1}}+\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}\tilde{A}h^{q}e^{hL}}}\right)\leq c_{1}\mathopen{}\mathclose{{\left(\frac{\tilde{A}}{B^{q}t^{q}}{}+\exp(-c_{2}t^{2})}}\right)$ (D.4)
with positive constants $c_{1}$ and $c_{2}$ only depending on $q$. We now adapt the idea of the original optimization argument by [Yurinsky1995, pp. 105] to the setting of our Lemma D.4. We first introduce the shorthand notation
$\Delta(t):=\tilde{A}/(B^{q}t^{q})\quad\text{and}\quad\Lambda(t):=L/(tB).$
With this notation, we rewrite the term on the left hand side of (D.4) as
$\frac{\Delta(t)}{\Lambda(t)^{q}}+\exp\mathopen{}\mathclose{{\left(-htB+htB\Delta(t)/\Lambda(t)^{q-1}+\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}(htB)^{q}\Delta(t)e^{\Lambda(t)thB}}}\right)$
$=\frac{\Delta(t)}{\Lambda(t)^{q}}+\exp\mathopen{}\mathclose{{\left(htB\mathopen{}\mathclose{{\left(\Delta(t)/\Lambda(t)^{q-1}-1}}\right)+\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}(htB)^{q}\Delta(t)e^{\Lambda(t)thB}}}\right).$ (D.5)
Note that for all $t>0$, $\Lambda(t)$ can be chosen independently of $\Delta(t)$ by adjusting the truncation level $L$ in our optimization. Note also that $\Delta(t)$ is exactly the first summand in the right hand side of the final desired bound (D.4). We will check the validity of the bound of the type (D.4) based on a case distinction.
Large $\Delta(t)$.
We assume $\Delta(t)\geq 1$. For every $t>0$, we may choose $L$ such that $\Lambda(t)>1$ and proceed to bound (D.5) as
$\frac{\Delta(t)}{\Lambda(t)^{q}}+\exp\mathopen{}\mathclose{{\left(htB\mathopen{}\mathclose{{\left(\Delta(t)/\Lambda(t)^{q-1}-1}}\right)+\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}(htB)^{q}\Delta(t)e^{\Lambda(t)thB}}}\right)$
$\leq\Delta(t)+\exp\mathopen{}\mathclose{{\left(htB\mathopen{}\mathclose{{\left(\Delta(t)-1}}\right)+\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}(htB)^{q}\Delta(t)e^{\Lambda(t)thB}}}\right)$
$\leq\Delta(t)+2\leq 3\Delta(t)+\exp(-c_{2}t^{2})$
for any $c_{2}>0$, if we choose $h$ small enough such that
$\exp(htB\underbrace{\mathopen{}\mathclose{{\left(\Delta(t)-1}}\right)}_{\geq 0}+\tilde{c}_{2}h^{2}B^{2}+\tilde{c}_{q}(htB)^{q}\Delta(t)e^{\Lambda(t)thB})\leq 2.$
This establishes (D.4).
Small $\Delta(t)$.
We therefore need to consider the case where $\Delta(t)<1$. Without loss of generality, in this setting we again assume that we are choosing $L$ appropriately to ensure that $\Lambda(t)>1$. This gives us
$\frac{\Delta(t)}{\Lambda(t)^{q}}\leq\Delta(t),$
and hence we only need to bound the exponential term in (D.5), as the polynomial term satisfies the final bound (D.4). We distinguish two cases, as the exponential term in (D.5) behaves differently for different ratios of $t^{2}/\log(1/\Delta(t))$.
Small $\Delta(t)$, case 1:
We assume $\Delta(t)<1$, $\Lambda(t)>1$ and $t^{2}\leq 6\tilde{c}_{2}\log(1/\Delta(t))$. We will show that in this case, the exponential term in (D.5) can be bounded by $c_{1}\exp(-c_{2}t^{2})$.
We choose the optimization parameter $h:=zt/B$ with some $z>0$, which we will determine later. The exponent in the right hand side of (D.5) can now be written as
$zt^{2}\mathopen{}\mathclose{{\left(\Delta(t)/\Lambda(t)^{q-1}\ -1}}\right)+z^{2}\tilde{c}_{2}t^{2}+\tilde{c}_{q}z^{q}t^{2q}\Delta(t)e^{\Lambda(t)zt^{2}}$
$=$ $t^{2}\mathopen{}\mathclose{{\left(\Delta(t)/\Lambda(t)^{q-1}-z+\tilde{c}_{2}z^{2}}}\right)+\tilde{c}_{q}z^{q}t^{2q}\Delta(t)e^{\Lambda(t)zt^{2}}$
$\leq$ $t^{2}\mathopen{}\mathclose{{\left(z\Delta(t)-z+\tilde{c}_{2}z^{2}}}\right)+\tilde{c}_{q}z^{q}(6\tilde{c}_{2}\log(1/\Delta(t))^{q}\Delta(t)^{1-6\tilde{c}_{2}\Lambda(t)z}$
$\leq$ $-\underline{c}_{1}t^{2}+6^{q}\tilde{c}_{q}\tilde{c}_{2}^{q}z^{q}e^{-q}\mathopen{}\mathclose{{\left(\frac{q}{\underline{c}_{2}}}}\right)^{q},$ (D.6)
where we invoke Lemma E.4 in the last step and choose $z$ small enough such that we have
$z\Delta(t)+z-\tilde{c}_{2}z^{2}>c_{1}>0\quad\text{and}\quad 1-6\tilde{c}_{2}\Lambda(t)z>\underline{c}_{2}>0,$
in which case the exponent estimate (D.6) is bounded by $c_{1}t^{2}+\tilde{c}$, yielding the claim.
Small $\Delta(t)$, case 2:
We assume $\Delta(t)<1$, $\Lambda(t)>1$ and $t^{2}>6\tilde{c}_{2}\log(1/\Delta(t))$. We will show that in this case, the exponential term in (D.5) can be bounded by $c_{1}\Delta(t)$.
In this regime, we choose the optimization parameter $h:=z\log(1/\Delta(t))/(tB)>0$ with some $z>0$ which we will determine later. The exponent in (D.5) with this choice becomes
$z\underbrace{\log(1/\Delta(t))}_{>0}(-1+\Delta(t)/\Lambda(t)^{q-1})+\tilde{c}_{2}z^{2}t^{-2}\underbrace{\log(1/\Delta(t))^{2}}_{<t^{2}\log(1/\Delta(t))/(6\tilde{c}_{2})}$
$+\tilde{c}_{q}z^{q}\log(1/\Delta(t))^{q}\Delta(t)e^{\Lambda(t)z\log(1/\Delta(t))}$
$\leq z\log(1/\Delta(t))(-1+\Delta(t))+\frac{z^{2}}{6}\log(1/\Delta(t))+\tilde{c}_{q}z^{q}\log(1/\Delta(t))^{q}\Delta(t)^{1-6\tilde{c}_{2}\Lambda(t)z}$
$=\underbrace{\log(\Delta(t))}_{<0}(z-\Delta(t)z-\frac{z^{2}}{6})+\tilde{c}_{q}z^{q}\log(1/\Delta(t))^{q}\Delta(t)^{1-6\tilde{c}_{2}\Lambda(t)z}$
$\leq\underbrace{\log(\Delta(t))}_{<0}+\tilde{c}_{q}z^{q}e^{-q}\mathopen{}\mathclose{{\left(\frac{q}{\underline{c}}}}\right)^{q},$ (D.7)
where in the last step we apply Lemma E.4 and choose $z$ small enough such that
$z-\Delta(t)z-\frac{z^{2}}{6}\geq 1\quad\text{and}\quad 1-6\tilde{c}_{2}\Lambda(t)z>\underline{c}>0.$
We finally obtain an upper bound of the exponent (D.7) as $\log(\Delta(t))+\tilde{c}$, yielding the claim.
Final bound:
The final bound (D.4) results from considering all of the the above cases and choosing the individual constants $c_{1}$ and $c_{2}$ large enough such that all cases are satisfied simultaneously (for example by choosing the maximum of the individual constants obtained above).
D.4 Concentration of empirical covariance operators